128 is a number was increased by 15% what was its original number?
Answer by fcabanski(1391)

Increased by 15% means multiply the number by (1 + 15/100).


Call the number x. x * (1 +.15) = 128


x + .15x = 1.15x = 128


x = 128/1.15 = 128/(115/100) = 128 * 100/115 = 12800/115